Work done on heating one mole of monoatomic gas adiabatically through 20°C is W. Then the work done on heating 6 moles of rigid diatomic gas through the same change in temperature
Ans 10w
1
Expert's answer
2020-10-23T11:53:14-0400

For the adiabatic process, from the first law of thermodynamics

"W = \Delta U"

The internal energy for one mole of ideal monoatomic gas is "U = \frac{3}{2}kT" .

The internal energy for one mole of ideal diatomic gas is "U = \frac{5}{2}kT" .

So, for the one mole of monoatomic gas "W_1 = \frac{3}{2} k \Delta T = 1.5 k \Delta T= W,"

For 6 moles of diatomic: "W_2 = 6 \cdot \frac{5}{2} k \Delta T = 15 k \Delta T"

"W_2 = 15 k \Delta T = 10 \cdot 1.5 k \Delta T = 10 W".


Answer: 10W.